# More is not always better? Enhancing Many-Shot In-Context Learning with Differentiated and Reweighting Objectives

### v1

@@ Abstract
Large language models (LLMs) excel at few-shot in-context learning (ICL) without requiring parameter updates. However, as the number of ICL demonstrations increases from a few to many, performance tends to plateau and eventually decline. We identify two primary causes for this trend: the suboptimal negative log-likelihood (NLL) optimization objective and the incremental data noise. To address these issues, we introduce DR-ICL, a novel optimization method that enhances model performance through Differentiated Learning and advantage-based Reweighting objectives. Globally, DR-ICL utilizes differentiated learning to optimize the NLL objective, ensuring that many-shot performance surpasses zero-shot levels. Locally, it dynamically adjusts the weighting of many-shot demonstrations by leveraging cumulative advantages inspired by reinforcement learning, thereby improving generalization. This approach allows the model to handle varying numbers of shots effectively, mitigating the impact of noisy data. Recognizing the lack of multi-task datasets with diverse many-shot distributions, we develop the Many-Shot ICL Benchmark (MICLB)—a large-scale benchmark covering shot numbers from 1 to 350 within sequences of up to 8,000 tokens—for fine-tuning purposes. MICLB facilitates the evaluation of many-shot ICL strategies across seven prominent NLP tasks and 50 distinct datasets. Experimental results demonstrate that LLMs enhanced with DR-ICL achieve significant improvements in many-shot setups across various tasks, including both in-domain and out-of-domain scenarios. We release the code and benchmark dataset hoping to facilitate further research in many-shot ICL [^1] https://github.com/xiaoqzhwhu/DR-ICL.
@@
Figure 1: The performance trend of LLMs across different $k$-shots scenarios. $k$ refers to the number of demonstration examples provided to LLMs, “+MetaICL” uses MetaICL for fine-tuning, while “+DR-ICL” uses our DR-ICL strategy.
@@ 1 Introduction
The performance trends of many-shot ICL can be attributed to several factors. The first factor is the training objective. As  Agarwal et al. highlights, while the straightforward NLL decreases during testing with ICL, performance on many downstream tasks also deteriorates. The second factor is the increasing noise with the large number of demonstrations.  Long et al.; Gao et al. demonstrate that the effectiveness of ICL heavily depends on the quality of the demonstrations. While in many-shot scenarios, utilizing a large number of high-quality demonstrations presents significant challenges, such as the huge workload of creating them and the difficulty of domain adaptation. Existing few-shot ICL methods do not address these issues, making them unsuitable for many-shot scenarios Zhang et al. (2024); Li et al. (2023); Agarwal et al. (2024); Bertsch et al. (2024).
@@
To address the above factors, we propose DR-ICL, enhancing many-shot in-context learning with a refined fine-tuning objective. For the first factor, we propose differentiated learning to deal with the trade-off between many-shot and zero-shot scenarios from a global perspective. During differentiated learning, we ensure that the performance on many-shot demonstrations surpasses that on zero-shot demonstrations. This approach promotes model’s deeper understanding of contextual cues, encouraging the model to effectively leverage contextual information. For the second factor, inspired by reinforcement learning, we propose an advantage-based reweighting method to filter noise in many-shot demonstrations from a local perspective. In reinforcement learning, the “advantage function” is essential for assessing the value of actions beyond the average expected return, effectively directing the policy to choose actions that are predicted to yield the highest future rewards Baird (1994). We divide each sequence containing $K$ demonstrations into multiple reweighting windows, each with $W$ demonstrations. For each demonstration $x_{k}$ in the reweighting window $w$, we designate the preceding window $w-1$ as the sampling window. We calculate the cumulative advantage from the data within the sampling window $w-1$ and apply it dynamically as the reward for $x_{k}$. Then, we integrate the cumulative advantage into the NLL computation for $x_{k}$ and get an advantage-based training objective. We combine the global differentiated learning and local reweighting as the final refined training objective for maintaining ICL under many-shots.
@@
In addition to the two challenges mentioned above, another significant obstacle is the lack of sufficient multi-task training data that spans a wide range of task numbers. Such data is crucial for studying the effects of in-context learning (ICL) across many-shot scenarios. We present MICLB, the largest dataset to study many-shot ICL, encompassing 50 tasks across 7 task types, and a total of over three million samples. The maximum number of shots achievable varies depending on the model, allowing for comprehensive investigations into many-shot ICL, including research on fine-tuning and inference. We categorize MICLB into different subsets based on task types, including in-domain and out-of-domain tasks, to fully assess the model’s ICL capabilities in many-shot scenarios.
@@
In summary, this paper identified two challenges in avoiding ICL’s decreasing performance under Many-shot scenarios: suboptimal training objective, and incremental data noise. We propose “differentiated learning” and “advantaged-based reweighting” to address these challenges, respectively. We further propose the largest MICLB to support our training, evaluation as well as further studies. We experiment DR-ICL on the MICLB with open-source LLMs, showing the stable performance both in-domain and out-of-domain under many-shots. All these points form the major contribution of this paper.
@@ 2 Related Work
Instruction Tuning of LLMs. Instruction tuning has become an effective technique for enhancing the capabilities and controllability of LLMs Zhang et al. (2023). In the domain of ICL, studies like MetaICL Min et al. (2022) and PEFT Bertsch et al. (2024) have demonstrated that fine-tuning LLMs with both small and large demonstration sizes, denoted as $k$, lead to improved ICL performance. Despite their studies being confined to a modest quantity of tuning data—capped at 10,000 entries—there is an evident necessity for deeper research into how ICL performs when scaled up with more extensive datasets. Consequently, we introduce MICLB, a significantly larger dataset, designed to delve into the strategies for amplifying ICL’s potential.
@@
Figure 2: The DR-ICL Training Framework. (a) The global differentiated learning for many-shot and zero-shot demonstrations. (b) The local advantage-based reweighting method assigns differential weights to demonstrations in window $w$ with window size $|W|=3$ and sampling size $|S|=1$, utilizing the cumulative advantage from the preceding window $w-1$.
3 DR-ICL
In this work, we propose the DR-ICL learning framework, which adjusts the weights of demonstrations and integrates reweighting within differentiated objectives, as illustrated in Figure 2. In DR-ICL, we organize training data through many-shot and zero-shot demonstrations. By simultaneously training the sequence of many-shot and zero-shot with a differentiated objective, we strengthen the model’s overall ICL capability. At the same time, to further reduce the noise of demonstrations in many-shot scenarios, we introduce a weighted training objective towards different samples in the many-shot demonstrations. By sampling the model’s performance under different demonstrations, we calculate the cumulative advantage gained as the number of demonstrations increases and use this cumulative advantage to adjust the learning process. Below, we show the components of the DR-ICL framework from both a global and local perspective, as well as the learning strategy.
@@ 3.2.1 Importance Sampling
Figure 3: The performance with incremental $k$-shots for Mistral-7B-Instruct-v0.2 and Llama-2-7b-chat-hf on CLSClusteringS2S under different strategies. We focus on CLSClusteringS2S for its high $k$-shot count, enabling a broader evaluation of DR-ICL. Our DR-ICL consistently shows better performance with a diverse range of $k$.
@@ 3.2.3 Reweighting
In the DR-ICL framework, we select important samples in the previous window and calculate the reward that measures the model’s performance in different positions to update the NLL loss for many-shot scenarios. We adjust the overall training objective of the many-shot sequence as follows:
@@ 3.2.4 Learning Strategy
The detailed process of the DR-ICL is presented in Algorithm 1. It enables the model to build upon prior knowledge at each iteration, avoiding uniform learning, thereby achieving progressive performance enhancement over long-term training.
Algorithm 1 Differentiated and Reweighting In-Context Learning (DR-ICL)
@@ 4.1.1 Datasets
To delve into the exploration of many-shot ICL in LLMs, we need plenty of data across a wide range of $k$-shots. The datasets employed in MetaICL, like CROSSFIT Ye et al. (2021) and UNIFIEDQA Khashabi et al. (2020), have a notable constraint: their task lengths are generally centered around 100 tokens. This focus restricts the wide range of $k$-shot distributions, especially when the training sequence length is constrained. On the other hand, the LongICLBench dataset introduced by Li et al. (2024) significantly extends the length ranging from 1,000 to 50,000 tokens. Nonetheless, the dataset’s limitation to a few hundred task instances renders it more suitable for inference rather than extensive training. In light of these limitations, we have developed the MICLB dataset. It encompasses 7 tasks of diverse difficulty levels and includes 50 datasets with average sample lengths per task that vary from 10 to 14,000 tokens. With the number of samples extending from the hundreds into the hundreds of thousands, the MICLB dataset ensures a substantial volume of data suitable for training and inference. More details can be found in the Appendix.
@@ 4.2 Results of Tasks
As shown in Tables 1, 2, 3, 5, and 6, DR-ICL significantly improves performance across various tasks. While MetaICL shows substantial fluctuations in $k$-shot performance on datasets like OpenbookQA and ARC, DR-ICL maintains more stable results. The slight advantage of our method over Meta-ICL is due to its focus on optimizing many-shot loss. IT’s performance declines with increasing context length, as it relies solely on zero-shot, which is less effective in many-shot scenarios. Additionally, Llama-2-7b-chat-hf’s 4,000-token limit causes performance on the ARC dataset to drop from 0.82 to 0.78 when $k$ exceeds 50. Under the DR-ICL framework, among the 12 datasets tested, $k=0$ led to a performance decrease on 5 datasets, no change on 2, and improvement on 5. Overall, performance improved by 0.5% with $k=0$, remaining stable. For $k>0$, datasets like CLSClusteringS2S showed continuous improvement, while DR-ICL effectively maintained performance stability as $k$ increased across most datasets.
@@ 4.3.1 Performance Tradeoff
We observe that the foundation models underperform on both datasets. After fine-tuning, the IT strategy achieves its best in the few-shot. MetaICL, benefiting from many-shot training data, performs well at larger $k$-shot levels but still shows significant fluctuations. In contrast, DR-ICL delivers more stable results, with accuracy steadily improving as $k$ increases. DR-ICL not only outperforms MetaICL in many-shot scenarios but also demonstrates faster loss convergence, as shown in Figure 6(a) in the Appendix 9, indicating its tradeoff of many-shot and zero-shot demonstrations.
@@ 4.3.2 Performance Variance
We track the performance variance across all datasets with NFT(6.49E-03), IT(2.71E-02), MetaICL(2.38E-03), and DR-ICL(1.56E-03) as $k$ varied. On average, DR-ICL has the lowest variance, suggesting a more stable performance with $k$-shot of demonstration changes. For details of the performance variance of each dataset please refer to Table 10 in Appendix 8.
@@ 4.3.3 Data Noise Sensitivity
We compare DR-ICL with and without local reweighting by examining how loss variance trends for each $k$-shot demonstration during training. The reweighting window in DR-ICL reduces loss variance and effectively balances the impact of noisy data by appropriately weighting demonstrations. This reduction in sensitivity to data noise helps maintain stable performance as the number of demonstrations increases. For details of the noise variation please refer to Table 11 in Appendix 11.
@@ 4.4.1 Hyperparameters
DR-ICL 0.30 0.51 0.5 0.53 0.55 0.57 0.48 0.63 0.52 0.51 0.63
DR-ICL w/ W=1 0.47 0.46 0.47 0.51 0.51 0.49 0.45 0.58 0.43 0.49 0.58
DR-ICL w/o global 0.45 0.43 0.46 0.41 0.52 0.51 0.52 0.50 0.43 0.47 0.52
DR-ICL w/o local 0.43 0.52 0.44 0.47 0.51 0.53 0.33 0.42 0.33 0.44 0.52
@@ 4.4.2 Global and Local Contribution
Table 7 displays the outcomes of DR-ICL when applying only global strategies or local strategies exclusively to the WinoWhy dataset. The results show that refining learning objectives via a global hyperparameter to trade off the performance and the local reweighting of demonstration examples can boost the LLMs’ ICL capabilities.
@@ 5 Conclusions
To enhance the ICL capacity as context lengths grow and demonstration $k$-shots rise, we introduce the DR-ICL algorithm to tackle the inaccurate objective and noise. This innovative method strategically calibrates global training goals to prioritize many-shot examples over zero-shot ones and employs local reweighting of many-shot instances using cumulative advantages as dynamic rewards, steering the model toward effective learning trajectories. To substantiate the effectiveness of our approach, we have curated and released the MICLB dataset, characterized by its diverse tasks and a broad spectrum of text lengths and quantities. Our method demonstrates notable enhancements in both in-domain and out-of-domain tasks. We anticipate that our research will stimulate further exploration into ICL’s potential and contribute to the advancement of LLM performance.
@@ B.2 Hyperparameters Study
Figure 6: (a) The many-shot training loss of DR-ICL converges to a lower level compared to IT and MetaICL. (b) The optimal performance is when the parameter $S$ is set to 1 on Llama-2-7b-chat-hf.
@@ B.3 Performance Variance
Table 10 is the performance variance across the NFT, IT, MetaICL, and DR-ICL methods. In Table 10, we prove that our method demonstrates the smallest deviation, indicating greater stability in performance as $k$-shot changes.
Dataset NFT IT MetaICL DR-ICL

### v2

@@ Abstract
Large language models (LLMs) excel at few-shot in-context learning (ICL) without requiring parameter updates. However, as the number of ICL demonstrations increases from a few to many, performance tends to plateau and eventually decline. We identify two primary causes for this trend: the suboptimal negative log-likelihood (NLL) optimization objective and the incremental data noise. To address these issues, we introduce DrICL, a novel optimization method that enhances model performance through Differentiated Learning and advantage-based Reweighting objectives. Globally, DrICL utilizes differentiated learning to optimize the NLL objective, ensuring that many-shot performance surpasses zero-shot levels. Locally, it dynamically adjusts the weighting of many-shot demonstrations by leveraging cumulative advantages inspired by reinforcement learning, thereby improving generalization. This approach allows the model to handle varying numbers of shots effectively, mitigating the impact of noisy data. Recognizing the lack of multi-task datasets with diverse many-shot distributions, we develop the Many-Shot ICL Benchmark (ICL-50)-a large-scale benchmark of 50 tasks that cover shot numbers from 1 to 350 within sequences of up to 8,000 tokens-for fine-tuning purposes. ICL-50 facilitates the evaluation of many-shot ICL strategies across seven prominent NLP tasks and 50 distinct datasets. Experimental results demonstrate that LLMs enhanced with DrICL achieve significant improvements in many-shot setups across various tasks, including both in-domain and out-of-domain scenarios. We release the code and benchmark dataset hoping to facilitate further research in many-shot ICL [^1] https://github.com/xiaoqzhwhu/DrICL.
@@
Figure 1: The performance trend of LLMs across different $k$-shots scenarios. $k$ refers to the number of demonstration examples provided to LLMs, “+MetaICL” uses MetaICL for fine-tuning, while “+DrICL” uses our DrICL strategy.
@@ 1 Introduction
We summarize two possible factors based on our preliminary study and previous works. The first factor is the training objective. As  Agarwal et al. highlights, while the straightforward NLL decreases during testing with ICL, performance on many downstream tasks also deteriorates. The second factor is the increasing noise with the large number of demonstrations.  Long et al.; Gao et al. demonstrate that the effectiveness of ICL heavily depends on the quality of the demonstrations. While in many-shot scenarios, utilizing a large number of high-quality demonstrations presents significant challenges, such as the huge workload of creating them and the difficulty of domain adaptation. Existing few-shot ICL methods do not address these issues, making them unsuitable for many-shot scenarios Zhang et al. (2024); Li et al. (2023); Agarwal et al. (2024); Bertsch et al. (2024).
@@
To address the above factors, we propose DrICL, enhancing many-shot in-context learning with a refined fine-tuning objective. For the first factor, we propose differentiated learning to deal with the trade-off between many-shot and zero-shot scenarios from a global perspective. During differentiated learning, we ensure that the performance on many-shot demonstrations surpasses that on zero-shot demonstrations. This approach promotes model’s deeper understanding of contextual cues, encouraging the model to effectively leverage contextual information. For the second factor, inspired by reinforcement learning, we propose an advantage-based reweighting method to filter noise in many-shot demonstrations from a local perspective. In reinforcement learning, the “advantage function” is essential for assessing the value of actions beyond the average expected return, effectively directing the policy to choose actions that are predicted to yield the highest future rewards Baird (1994). We divide each sequence containing $K$ demonstrations into multiple reweighting windows, each with $W$ demonstrations. For each demonstration $x_{k}$ in the reweighting window $w$, we designate the preceding window $w-1$ as the sampling window. We calculate the cumulative advantage from the data within the sampling window $w-1$ and apply it dynamically as the reward for $x_{k}$. Then, we integrate the cumulative advantage into the NLL computation for $x_{k}$ and get an advantage-based training objective. We combine the global differentiated learning and local reweighting as the final refined training objective for maintaining ICL under many-shots.
@@
In addition to the two challenges mentioned above, another significant obstacle is the lack of sufficient multi-task training data that spans a wide range of task numbers. Such data is crucial for studying the effects of ICL across many-shot scenarios. We present ICL-50, the largest dataset to study many-shot ICL, encompassing 50 tasks across 7 task types, and a total of over three million samples. The maximum number of shots achievable varies depending on the model, allowing for comprehensive investigations into many-shot ICL, including research on fine-tuning and inference. We categorize ICL-50 into different subsets based on task types, including in-domain and out-of-domain tasks, to fully assess the model’s ICL capabilities in many-shot scenarios.
@@
In summary, this paper identified two challenges in avoiding ICL’s decreasing performance under Many-shot scenarios: suboptimal training objective, and incremental data noise. We propose “differentiated learning” and “advantaged-based reweighting” to address these challenges, respectively. We further propose the largest ICL-50 to support our training, evaluation as well as further studies. We experiment DrICL on the ICL-50 with open-source LLMs, showing the stable performance both in-domain and out-of-domain under many-shots. All these points form the major contribution of this paper.
@@ 2 Related Work
Instruction Tuning of LLMs. Instruction tuning has become an effective technique for enhancing the capabilities and controllability of LLMs Zhang et al. (2023). In the domain of ICL, studies like MetaICL Min et al. (2022) and PEFT Bertsch et al. (2024) have demonstrated that fine-tuning LLMs with both small and large demonstration sizes, denoted as $k$, lead to improved ICL performance. Despite their studies being confined to a modest quantity of tuning data—capped at 10,000 entries—there is an evident necessity for deeper research into how ICL performs when scaled up with more extensive datasets. Consequently, we introduce ICL-50, a significantly larger dataset, designed to delve into the strategies for amplifying ICL’s potential.
@@
Figure 2: The DrICL Training Framework. (a) The global differentiated learning for many-shot and zero-shot demonstrations. (b) The local advantage-based reweighting method assigns differential weights to demonstrations in window $w$ with window size $|W|=3$ and sampling size $|S|=1$, utilizing the cumulative advantage from the preceding window $w-1$.
3 DrICL
In this work, we propose the DrICL learning framework, which adjusts the weights of demonstrations and integrates reweighting within differentiated objectives, as illustrated in Figure 2. In DrICL, we organize training data through many-shot and zero-shot demonstrations. By simultaneously training the sequence of many-shot and zero-shot with a differentiated objective, we strengthen the model’s overall ICL capability. At the same time, to further reduce the noise of demonstrations in many-shot scenarios, we introduce a weighted training objective towards different samples in the many-shot demonstrations. By sampling the model’s performance under different demonstrations, we calculate the cumulative advantage gained as the number of demonstrations increases and use this cumulative advantage to adjust the learning process. Below, we show the components of the DrICL framework from both a global and local perspective, as well as the learning strategy.
@@ 3.2.1 Importance Sampling
Figure 3: The performance with incremental $k$-shots for Mistral-7B-Instruct-v0.2 and Llama-2-7b-chat-hf on CLSClusteringS2S under different strategies. We focus on CLSClusteringS2S for its high $k$-shot count, enabling a broader evaluation of DrICL. Our DrICL consistently shows better performance with a diverse range of $k$.
@@ 3.2.3 Reweighting
In the DrICL framework, we select important samples in the previous window and calculate the reward that measures the model’s performance in different positions to update the NLL loss for many-shot scenarios. We adjust the overall training objective of the many-shot sequence as follows:
@@ 3.2.4 Learning Strategy
The detailed process of the DrICL is presented in Algorithm 1. It enables the model to build upon prior knowledge at each iteration, avoiding uniform learning, thereby achieving progressive performance enhancement over long-term training.
Algorithm 1 Differentiated and Reweighting In-Context Learning (DrICL)
@@ 4.1.1 Datasets
To delve into the exploration of many-shot ICL in LLMs, we need plenty of data across a wide range of $k$-shots. The datasets employed in MetaICL, like CROSSFIT Ye et al. (2021) and UNIFIEDQA Khashabi et al. (2020), have a notable constraint: their task lengths are generally centered around 100 tokens. This focus restricts the wide range of $k$-shot distributions, especially when the training sequence length is constrained. On the other hand, the LongICLBench dataset introduced by Li et al. (2024) significantly extends the length ranging from 1,000 to 50,000 tokens. Nonetheless, the dataset’s limitation to a few hundred task instances renders it more suitable for inference rather than extensive training. In light of these limitations, we have developed the ICL-50 dataset. It encompasses 7 tasks of diverse difficulty levels and includes 50 datasets with average sample lengths per task that vary from 10 to 14,000 tokens. With the number of samples extending from the hundreds into the hundreds of thousands, the ICL-50 dataset ensures a substantial volume of data suitable for training and inference. More details can be found in the Appendix.
@@ 4.2 Results of Tasks
As shown in Tables 1, 2, 3, 5, and 6, DrICL significantly improves performance across various tasks. While MetaICL shows substantial fluctuations in $k$-shot performance on datasets like OpenbookQA and ARC, DrICL maintains more stable results. The slight advantage of our method over Meta-ICL is due to its focus on optimizing many-shot loss. IT’s performance declines with increasing context length, as it relies solely on zero-shot, which is less effective in many-shot scenarios. Additionally, Llama-2-7b-chat-hf’s 4,000-token limit causes performance on the ARC dataset to drop from 0.82 to 0.78 when $k$ exceeds 50. Under the DrICL framework, among the 12 datasets tested, $k=0$ led to a performance decrease on 5 datasets, no change on 2, and improvement on 5. Overall, performance improved by 0.5% with $k=0$, remaining stable. For $k>0$, datasets like CLSClusteringS2S showed continuous improvement, while DrICL effectively maintained performance stability as $k$ increased across most datasets.
@@ 4.3.1 Performance Tradeoff
We observe that the foundation models underperform on both datasets. After fine-tuning, the IT strategy achieves its best in the few-shot. MetaICL, benefiting from many-shot training data, performs well at larger $k$-shot levels but still shows significant fluctuations. In contrast, DrICL delivers more stable results, with accuracy steadily improving as $k$ increases. DrICL not only outperforms MetaICL in many-shot scenarios but also demonstrates faster loss convergence, as shown in Figure 6(a) in the Appendix 9, indicating its tradeoff of many-shot and zero-shot demonstrations.
@@ 4.3.2 Performance Variance
We track the performance variance across all datasets with NFT(6.49E-03), IT(2.71E-02), MetaICL(2.38E-03), and DrICL(1.56E-03) as $k$ varied. On average, DrICL has the lowest variance, suggesting a more stable performance with $k$-shot of demonstration changes. For details of the performance variance of each dataset please refer to Table 10 in Appendix 8.
@@ 4.3.3 Data Noise Sensitivity
We compare DrICL with and without local reweighting by examining how loss variance trends for each $k$-shot demonstration during training. The reweighting window in DrICL reduces loss variance and effectively balances the impact of noisy data by appropriately weighting demonstrations. This reduction in sensitivity to data noise helps maintain stable performance as the number of demonstrations increases. For details of the noise variation please refer to Table 11 in Appendix 11.
@@ 4.4.1 Hyperparameters
DrICL 0.30 0.51 0.5 0.53 0.55 0.57 0.48 0.63 0.52 0.51 0.63
DrICL w/ W=1 0.47 0.46 0.47 0.51 0.51 0.49 0.45 0.58 0.43 0.49 0.58
DrICL w/o global 0.45 0.43 0.46 0.41 0.52 0.51 0.52 0.50 0.43 0.47 0.52
DrICL w/o local 0.43 0.52 0.44 0.47 0.51 0.53 0.33 0.42 0.33 0.44 0.52
@@ 4.4.2 Global and Local Contribution
Table 7 displays the outcomes of DrICL when applying only global strategies or local strategies exclusively to the WinoWhy dataset. The results show that refining learning objectives via a global hyperparameter to trade off the performance and the local reweighting of demonstration examples can boost the LLMs’ ICL capabilities.
@@ 5 Conclusions
To enhance the ICL capacity as context lengths grow and demonstration $k$-shots rise, we introduce the DrICL algorithm to tackle the inaccurate objective and noise. This innovative method strategically calibrates global training goals to prioritize many-shot examples over zero-shot ones and employs local reweighting of many-shot instances using cumulative advantages as dynamic rewards, steering the model toward effective learning trajectories. To substantiate the effectiveness of our approach, we have curated and released the ICL-50 dataset, characterized by its diverse tasks and a broad spectrum of text lengths and quantities. Our method demonstrates notable enhancements in both in-domain and out-of-domain tasks. We anticipate that our research will stimulate further exploration into ICL’s potential and contribute to the advancement of LLM performance.
@@ B.2 Hyperparameters Study
Figure 6: (a) The many-shot training loss of DrICL converges to a lower level compared to IT and MetaICL. (b) The optimal performance is when the parameter $S$ is set to 1 on Llama-2-7b-chat-hf.
@@ B.3 Performance Variance
Table 10 is the performance variance across the NFT, IT, MetaICL, and DrICL methods. In Table 10, we prove that our method demonstrates the smallest deviation, indicating greater stability in performance as $k$-shot changes.
Dataset NFT IT MetaICL DrICL